### v1

# Noise-resilient host for electron qubit operation up to 0.4 K

Noise-resilient host for electron qubit operation up to 0.4 K
Xinhao Li Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA Department of Physics, Harvard University, Cambridge, Massachusetts 02138, USA    Christopher S. Wang James Franck Institute and Department of Physics, University of Chicago, Chicago, Illinois 60637, USA    Brennan Dizdar James Franck Institute and Department of Physics, University of Chicago, Chicago, Illinois 60637, USA    Yizhong Huang Pritzker School of Molecular Engineering, University of Chicago, Chicago, Illinois 60637, USA Department of Electrical and Computer Engineering, Northeastern University, Boston, Massachusetts 02115, USA    Yutian Wen Department of Physics and Astronomy, University of Notre Dame, Notre Dame, Indiana 46556, USA    Wei Guo National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA Department of Mechanical Engineering, FAMU-FSU College of Engineering, Florida State University, Tallahassee, Florida 32310, USA    Xufeng Zhang Department of Electrical and Computer Engineering, Northeastern University, Boston, Massachusetts 02115, USA    Xu Han xu.han@anl.gov Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA Pritzker School of Molecular Engineering, University of Chicago, Chicago, Illinois 60637, USA    Xianjing Zhou xianjing.zhou@cea.fr National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA Department of Mechanical Engineering, FAMU-FSU College of Engineering, Florida State University, Tallahassee, Florida 32310, USA    Dafei Jin dfjin@nd.edu Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA Department of Physics and Astronomy, University of Notre Dame, Notre Dame, Indiana 46556, USA
(February 3, 2025)
Cryogenic solid neon has recently emerged as a pristine solid-state material platform for hosting single electron qubits. At temperatures of $\sim$10  mK, electron-on-solid-neon (eNe) charge qubits have shown exceptionally long coherence times and high single-qubit gate fidelities. Further expanding the capabilities of this platform to multiple qubits in a scalable fashion will require careful understanding and the corresponding optimization of the underlying noise environment. Here, we demonstrate the remarkable robustness enabled by the neon layer in eNe charge qubits against charge and thermal noise by operating them away from the charge-insensitive sweet-spot and at elevated temperatures. Without optimization, the measured charge (voltage) noise spectral density on neon is comparable to the best in conventional gate-defined semiconductor quantum dots. Up to 0.4 K, the qubit coherence time with dynamical decoupling can still reach the relaxation limit of $\sim$1 microsecond. These observations highlight the potential of solid neon to facilitate a quiet charge environment and eNe qubits’ potential for higher-temperature operation in a scaled-up system.
Solid-state electron qubits are inherently affected by decoherence mechanisms in their host materials. Spectral characterization has revealed a typical $1/f$ noise distribution for both charge and spin qubits, often attributed to individual charge fluctuators at the surfaces and interfaces of semiconductor or superconducting materials [1, 2, 3]. Extensive effort has been made to extend electron qubit coherence times by reducing the noise density in the surrounding environment and minimizing the sensitivity of qubits to noise  [4, 2, 5]. Recently, we demonstrated solid neon as a novel host material that traps electrons at the neon-vacuum interface [6, 7]. Using a cavity quantum electrodynamics architecture, we address the charge states of eNe by coupling them to superconducting resonators. At the charge sweet-spot, eNe qubits are first-order insensitive to charge noise, showing long coherence times $T_{2}^{*}$ of up to $\sim$50 $\mu$s (ref [7]), which is nearly four orders of magnitude longer than for reported semiconductor charge qubits [8]. In principle, this should translate quite favorably for operating eNe spin qubits, assuming that isotopically purified neon is used [9, 10, 11].
There remains, however, more to be understood about the nature of eNe charge qubits (which we will simply refer to as “eNe qubits” for the remainder of the manuscript) and their environment. The specific electron trapping mechanism on solid neon is unclear, though it is likely that disorder in the neon surface plays a key role [12]. Moreover, the environmental coherence-limiting factors in current devices remain elusive. Investigating the performance of eNe qubits away from the sweet-spot and at elevated temperatures, when the qubit is subject to charge and thermal noise, can provide valuable insight on the coupling between eNe and the environment, revealing the role of the neon host and paving the way for improved qubit performance and stability. This is crucial for scaling up the eNe qubit platform, as environmental noise acts as a major obstacle to the precise creation and consistent retention of entangled multi-qubit states [1, 13]. It is also important for understanding the limitations of spin-state control [2, 14, 9] of eNe using electrically sensitive mechanisms such as synthetic spin-orbit coupling [15, 16, 17, 18, 19] or exchange interactions [20, 21, 22, 23]. Furthermore, operating qubits at elevated temperatures can mitigate many engineering constraints due to limited cooling power at base temperature, advantageous for scaling [24, 25, 26].
Fig. 1. : eNe charge qubit coupled to a TiN high-impedance superconducting resonator. a, Illustration of the high impedance TiN superconducting resonator with two identical electron traps. Microwave (MW) signals to control and read out the electron qubit’s charge state are sent through the resonator via the input and output ports. Gate electrodes are patterned with on-chip low-pass filters to apply DC bias voltage to the qubits while maintaining the resonator’s quality factor. The metal plane between the resonator pins is connected to the ground plane via aluminum wire bonds. b, False-color scanning electron micrograph image of the electron trapping area on the right end of the resonator with trap gates (RTG). c, Cross-section schematic of the electron trapping area.
[…]
Fig. 2. : Spectroscopic and coherence properties of an eNe charge qubit. a, Normalized microwave transmission amplitude ($A/A_{0}$) centered around the resonator frequency versus the relative resonator bias voltage $\Delta V_{\text{res}}$ as described in (b). Two avoided crossings appear when the eNe qubit comes into resonance with the resonator. b, Two-tone qubit spectroscopy measurement displaying the transmission phase response at the resonator frequency $\omega_{\text{r}}$ versus $\Delta V_{\text{res}}$, for different weak probe tones concurrently sent in at $f_{\text{drive}}$. c, Ramsey fringes at the charge sweet-spot, marked with the red arrow in (b), with fitted $T_{2}^{*}$ of 8.2 µs. $P_{e}$ is the qubit’s excited-state population. d, $T_{1}$ = 11.6 µs and $T_{2}^{\text{echo}}$ = 21.6 µs at the charge sweet-spot.
Device structure
Our device consists of a split superconducting resonator made of a 30 nm thick TiN film grown on a $\langle 111\rangle$-oriented intrinsic silicon (Si) substrate by atomic layer deposition [27], as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka. An electron trap is positioned at each end of the resonator, with the broader goal of coupling two distant eNe qubits via the resonator bus [28]. Each electron trap features a simplified rectangular structure compared to the previously used oval-shaped design [6, 7]. Around the traps, the Si substrate is etched down by approximately 250 nm to host the thin neon layer, as illustrated in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb. In this work, we utilize the differential mode of the resonator to couple to motional states of electrons trapped on solid neon (see Supplementary I), with the microwave electrical field pointing from one resonator pin to the other [29], as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kc. Considering the first two charge states of eNe, the coupled system can be described by the Jaynes-Cummings Hamiltonian [14]:
$H=\hbar\omega_{\text{r}}\left(a^{\dagger}a+\frac{1}{2}\right)+\frac{1}{2}\hbar\omega_{\text{q}}\sigma_{z}+g(a^{{\dagger}}\sigma_{-}+a\sigma_{+}),$ (1)
where $\omega_{\text{r}}/2\pi=5.668$ GHz is the resonator frequency after neon deposition, $\omega_{\text{q}}/2\pi$ is the qubit transition frequency, $g$ is the electron-photon coupling strength, $a^{{\dagger}}$ and $a$ are the photon creation and annihilation operators, respectively, $\sigma_{z}$ and $\sigma_{\pm}\equiv\sigma_{x}\pm i\sigma_{y}$ are the standard Pauli operators on a two-level system.
As an improvement from previous work, we leverage the high kinetic inductance ($\sim$20 pH/$\Box$) of the thin TiN film to enhance the qubit-resonator coupling strength [30]. The estimated equivalent lumped element impedance for the differential mode is $Z_{r}\sim 600\,\Omega$ (see Supplementary I), approximately ten times that of the previous niobium (Nb) device [6, 29]. Since $g\propto\omega_{r}\sqrt{Z_{r}}$, we expect the high impedance resonator to support a coupling strength achieving $g/2\pi\sim 10\,\text{MHz}$ level [30]. Meanwhile, several DC gates are placed around the electron trapping area to tune the qubit transition frequency, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka-b. In order to minimize microwave leakage of the resonator mode through the DC gates, all gates are equipped with on-chip low-pass filters with a $\sim$0.5 GHz cutoff frequency providing over 60 dB attenuation at resonator frequency. This is particularly important in our device given that for high-impedance resonators, the parasitic gate capacitance can be comparable to the resonator capacitance [31, 32]. With this design, the resonator maintains a narrow linewidth $\kappa/2\pi$ of 0.38 MHz (see Supplementary I). Before performing electron experiments, a thin layer of neon was grown on the surface of the device following the same procedure as in our previous work [7], resulting in less than 1 MHz redshift of the resonator frequency. Electrons are emitted from a tungsten filament and bound to the surface of the neon film. Only when both the neon film is present on the sample surface and electrons are emitted do we see signatures of trapped eNe qubits strongly coupled to the resonator (see Supplementary II).
Fig. 3. : Dephasing of eNe charge qubit Q1. a, Calculated $1/T_{e}^{\rm{CPMG}}$ for various qubit bias points and refocusing pulse numbers $N$. With the increase of $N$, the $1/T_{e}^{\rm{CPMG}}$ of the qubit biased near the charge sweet-spot approaches the limit of $2T_{1}$. b, The pure dephasing time $T_{\phi}$ increases as a function of $N$ when the qubit biased away from charge sweet-spot, with a power-law fit of $T_{\phi}\propto N^{0.61}$. c, Repeated Ramsey fringes measured near qubit sweet-spot for 128 iterations, with each record taking 33 s. d, Detuning $\Delta_{dq}$ between drive tone and qubit frequency during the Ramsey measurements, revealing stochastic frequency shifts.
Qubit performance
As illustrated in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kc, the local neon profile and the trap structure define the potential energy landscape seen by the electrons. Recent theoretical work emphasized the important role that the neon surface plays in defining the qubit’s Hamiltonian [12], whose exact form remains unclear. Here we approximate each eNe qubit’s transition frequency with a generalized hyperbolic model, capturing the measured qubit spectroscopic features:
$\hbar\omega_{\text{q}}\,=\,\sqrt{h^{2}(\omega_{\text{ss}}+\delta\omega_{\text{ss}})^{2}+(\epsilon+\delta\epsilon+2d_{e}(\mathcal{E}+\delta\mathcal{E}))^{2}},$ (2)
where $\omega_{\text{ss}}$ represents the the charge sweet-spot frequency, and $\epsilon$ describes the energy off-set defining the corresponding bias voltage. This model has been widely applied to describe two-level quantum systems [33], with an energy landscape consistent with our bounds of large anharmonicity in eNe charge qubits [7]. The electrical tunability of the qubit’s transition frequency is described by the term $2d_{\rm{e}}\mathcal{E}$, where $d_{\rm{e}}$ is the electron dipole moment and $\mathcal{E}$ is the applied field [34, 35, 36]. Meanwhile, noise terms $\delta\mathcal{E}$, $\delta\epsilon$ and $\delta\omega_{\text{ss}}$ caused by DC bias or adjacent charge fluctuations lead to qubit decoherence.
Bringing eNe qubit’s transition frequency $\omega_{\rm{q}}$ onto resonance with $\omega_{\rm{r}}$ results in the vacuum Rabi splitting in the resonator’s transmission spectrum. An example of this is shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka for Q1, one of the three qubits we characterized, with the electron-photon coupling strength $g/2\pi=6.43$ MHz and on-resonance qubit linewidth $\gamma/2\pi=3.81$ MHz (see Supplementary III-VI for details of characterizations on the three qubits Q1-3). Compared to the previous Nb resonator [6, 7], the higher impedance of the TiN resonator enhances the qubit-resonator coupling strength, with a maximum observed $g/2\pi$ of approximately $16$ MHz (see Q3 in Supplementary VI). We further mapped the qubit spectrum by applying a second drive tone and probing at the resonator frequency at low power, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb. The extracted qubit frequency follows a hyperbolic dependence on the DC voltage $V_{\text{res}}$ applied to the resonator with a charge sweet-spot at 5.065 GHz (see Supplementary IV). The sweet-spot frequency, coupling strength, and sensitivity to DC biases ($\partial f_{\text{q}}/\partial V_{\text{res}}$) vary between qubits (see Supplementary III). This variation suggests some randomness in the local trapping potential, which determines the qubit’s minimum transition frequency and the projection of the electron’s dipole moment along the resonator mode and the applied DC field.
Next, we characterize the coherence performance of the qubit biased at its charge sweet-spot. Figure Noise-resilient host for electron qubit operation up to 0.4 Kc-d show the measured $T_{1}$, $T_{2}^{*}$, and $T_{2}^{\text{echo}}$ of 11.6 µs, 8.2 µs, and 22.6 µs, respectively. Unlike the qubit reported in our previous works [7], the relaxation time ($T_{1}$) of eNe qubits observed on the new TiN device is generally not Purcell-limited at their charge sweet-spots, with over 100 MHz detuning from the resonator ($\Delta_{\rm{rq}}/2\pi$). The Purcell rate of Q1 to resonator mode is $\Gamma_{\rm{r}}\,=\,\kappa g^{2}/\Delta_{\rm{rq}}^{2}\,=\,1/3.9\,\text{ms}^{-1}$, which indicates that non-radiative decay channels dominate the energy relaxation of Q1. Measurements of the Q1’s $T_{1}$ at various bias frequencies (see Noise Spectroscopy section) further confirm that the non-radiative decay dominates the energy relaxation unless the qubit frequency is tuned much closer to the resonator frequency. Additionally, the fact that the $T_{2}^{\rm{echo}}$ approaches $2T_{1}$ indicates that the quasi-static noise is the dominating dephasing factor for Q1 at its charge sweet-spot [37, 38].
Fig. 4. : Noise spectroscopy of qubit Q1. Main plot between 0.01 $\sim$ 1 MHz: Colored dots: Total noise density derived from dynamical decoupling data at different qubit bias points. Main plot between $10^{-3}$ to $10^{-1}$ Hz: Blue dots: Calculated total noise density from long-term Ramsey measurements when biased at charge sweet-spot. Main plot near 5 GHz: Green triangles: Transverse noise of the eNe qubit. Inset plot i: Equivalent charge (voltage) noise on the resonator electrode calculated. Inset plot ii: Zoom-in of transverse noise. Gray dashed lines: power-law fits of frequency dependent noise.
Noise spectroscopy
Fluctuations in the charge environment can cause stochastic frequency shifts of the eNe qubit, which leads to qubit dephasing. Armed with high fidelity control and a known voltage (charge) lever arm, we first characterize the high-frequency noise spectral density of Q1 using the Carr-Purcell-Meiboom-Gill (CPMG) sequence with $N=0,1,2,4,6,8,12$, and 16 refocusing pulses. All bias points had positive $\Delta V_{\rm{res}}$ as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb, with a maximum sensitivity of $\partial f_{\rm{q}}/\partial V_{\rm{res}}=180.7$ MHz/mV, corresponding to a qubit frequency detune of 15.9 MHz from the charge sweet-spot. The measured average energy relaxation time at all bias points is 11.94 $\pm$ 0.3 µs, indicating that the same non-radiative decay channel dominates. We fit the measured coherence decay to extract the decoherence time $T_{2}^{\text{CPMG}}$ and pure dephasing time $T_{\phi}$ (See Methods). As shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka, in the absence of refocusing pulses (i.e., the Ramsey measurement), the decoherence rate $1/T_{2}^{\text{CPMG}}$ increases with sensitivity to charge noise. At $\partial f_{\rm{q}}/\partial V_{\rm{res}}=180.7$ MHz/mV, $T_{2}^{*}$ decreases to 1.93 µs. Introducing refocusing pulses enhances $T^{\text{CPMG}}_{2}$ at every biasing point, approaching two times of $T_{1}$ as $N$ increased. At the charge sweet-spot, the rapid saturation of $T_{2}^{\text{CPMG}}$ with increasing $N$ indicates a low noise power density in the relatively high frequency range (100 kHz). When the qubit was biased at a point more sensitive to charge noise, more pulses were required to extend the $T_{2}^{\text{CPMG}}$ towards the energy relaxation limit. In Supplementary V and VI, we show that for the other two eNe qubits, a single refocusing pulse is insufficient to mitigate the majority of noises at the qubits’ sweet-spot, suggesting higher noise density at the high-frequency range. This variation in noise behavior suggests some complexities of the local noise environment experienced by individual electrons on the neon surface.
[…]
Material platform Qubit type $S_{v}$ ($\mu\text{V}^{2}/$Hz) Noise frequency (kHz) $\partial f_{q}/\partial V$ (MHz/mV) Reference Neon single-electron charge $10^{-5}\sim 10^{-7}$ $10\sim 1000$ $10\sim$100 this work Si/SiO2 singlet-triplet spin $10^{-1}\sim 10^{-2}$ $100\sim 1000$ $\sim$ 10 ref. [20] Si/SiO2 single-electron spin $10^{-1}\sim 10^{-2}$ $10\sim 100$ $0.2$ ref. [15] 28Si/28SiO2 single-electron spin $10^{0}\sim 10^{-1}$ $1\sim 10$ $0.01$ ref. [16] Si/SiGe singlet-triplet spin $10^{-4}\sim 10^{-5}$ $100\sim 1000$ $\sim 10$ ref. [21] Si/SiGe single-electron spin $10^{-1}\sim 10^{-3}$ $10\sim 1000$ $0.55$ ref. [19] 28Si/SiGe single-electron spin $10^{-2}\sim 10^{-3}$ $10\sim 100$ $0.09$ ref. [17] 28Si/SiGe single-electron spin $10^{-1}$ $10$ $0.36$ ref. [18] 28Si/SiGe singlet-triplet spin $10^{-2}\sim 10^{-4}$ $10\sim 1000$ $4.5$ ref. [22] GaAs/AlGaAs singlet-triplet spin $10^{-7}\sim 10^{-8}$ $100\sim 1000$ $10\sim 100$ ref. [23]
Tab. 1. : Comparison between voltage noises experienced by electron qubits on solid neon and in semiconductor material platforms.
To benchmark different electron qubit material platforms, we compare the environmental charge noise measured as voltage fluctuations on adjacent gate electrodes, situated at a distance on the scale of around one hundred nanometers from the qubits. Specifically, we examine electrons trapped on solid neon and those confined in semiconductor materials, which are the leading platforms for hosting electron qubits. Table 1 summarizes our results on eNe and compares them with data from literature on semiconductor qubits. Because of the clean qubit environment provided by the neon interface, the charge noise experienced by eNe is orders of magnitude lower than that in typical semiconductor materials, approaching some of the best performance. Since the susceptibility of electron spin qubit to charge noise through effects including spin-orbit coupling and exchange interaction is in general less than that of electron charge qubit, as shown in the “$\partial f_{q}/\partial V$” column of Tab. Noise-resilient host for electron qubit operation up to 0.4 K, we foresee that electron spin qubit on solid neon could have even better coherence. Further development of the electron trap and the suppression of the influence from redundant electrons on neon will further improve the performance of eNe qubits.
After investigating high-frequency noise, we turn to the low-frequency noise at the charge sweet-spot. Figure Noise-resilient host for electron qubit operation up to 0.4 Kc-d show variations in Q1’s frequency revealed by performing repeated Ramsey measurements for approximately one hour [21, 20] since initially biased near its charge sweet-spot. The qubit undergoes a discrete frequency transition of varying magnitudes, as seen near the 25 minutes of the one-hour tracing, revealed by a Fast Fourier transform (FFT) of the Ramsey signals (See Methods and Supplementary IV). This suggests slow mechanisms driving qubit decoherence, akin to low-frequency dynamics in semiconductor and superconducting qubits [21, 20, 40], which can be mitigated with feedback control [41, 42]. Using a periodogram method, we convert this one-hour qubit frequency measurement into a frequency noise spectrum over $10^{-3}$ to $10^{-1}$ Hz, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 K. A power-law fit of the data gives the relation of $S\propto f^{-1.11}$. Despite the qubit is first-order insensitive to small charge (voltage) fluctuation at the sweet-spot, it can still experience slow frequency drifts due to nearby charge fluctuations, as observed in both semiconductor and superconducting qubits via second-order effects [43, 44]. Other detection methods, such as single-electron charge sensing techniques [21], could complete the noise spectrum ranging from 1 Hz to $10^{4}$ Hz.
To complete the discussion, we calculate the transverse noise at the qubit frequencies contributing to energy relaxation with $1/T_{1}=\pi/2\times S(2\pi f_{\rm{q}})$ (ref. [37]). The fitted total transverse noise versus frequency follows $S\propto f^{2.22}$ between 5.065 to 5.498 GHz, with a mean of $3.6\times 10^{5}$$^{2}/\text{Hz}$ Hz, as shown in the inset of Fig. Noise-resilient host for electron qubit operation up to 0.4 K.
Temperature dependence
To obtain more information about the noise environment of eNe qubits, we measure the temperature dependence of Q1’s coherence ($T_{1}$, $T_{2}^{*}$, and $T_{2}^{\rm{echo}}$) at its charge sweet-spot from 10 mK to 500 mK. The energy relaxation $T_{1}$ is well described by a model that only accounts for the coupling of a two-level quantum system to a bosonic thermal bath within the experiment’s temperature range, $T_{1}\propto\text{tanh}(\hbar\omega_{\rm{q}}/2k_{\text{B}}T)$ (Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka) (ref. [45, 46]). Similarly, the thermal population follows that of a Maxwell-Boltzmann distribution [47] (See Supplementary IV). The corresponding single electron temperature closely tracks the mixing-chamber (MXC) temperature between 100 to 500 mK and saturates around 40 mK, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Ka. This effective cooling of electron qubits below 100 mK is difficult to achieve with electrons in semiconductors, due to the suppressed electron-environment interaction at low temperatures [48].
A separate measurement on a different qubit Q2, presented in Supplementary V, also reveals strikingly thermal behavior in $T_{1}$ and thermal population but with a different $T_{1}$ in the low-temperature limit. It suggests a picture in which individual eNe qubits couple differently to their environment (perhaps through phonons [14, 9]) which sets the low temperature $T_{1}$. This motivates further studies of the microscopic limitations of $T_{1}$. At these qubit frequencies, $T_{1}$ drops to only about half the low-temperature value at 200 mK, facilitating higher temperature operation. In contrast, electron qubits in semiconductor materials are susceptible to phonon effects [25, 49], which can significantly degrade $T_{1}$ at elevated temperatures. Given these favorable scaling, we can also anticipate further improvement of the high-temperature performance of eNe qubits by engineering the electron trapping and detection structure to support higher qubit frequency operation [24].
The temperature dependence of the coherence data, however, remains complex. $T_{2}^{*}$ and $T_{2}^{\rm{echo}}$ show that quasi-static and high-frequency noise components behave differently with temperature, as shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kb. With increasing temperature, $T_{2}^{*}\propto T^{-0.74}$, indicating the quasi-static noise power scales almost linearly against temperature. Below 50 mK, $T_{2}^{\rm{echo}}$ approaches $2T_{1}$, suggesting that the single echo pulse still effectively mitigates the quasi-static noise that contributes to most of the dephasing. However, starting from 75 mK, $T_{2}^{\rm{echo}}$ scales with temperature as $\propto T^{-1.1}$, and quickly degrades from $2T_{1}$, indicating the rise of non-quasistatic dephasing noises at higher temperatures [23]. Despite that, Q1 maintains $>$ 1 µs $T_{2}^{\rm{echo}}$ at elevated temperatures up to 400 mK, showing eNe’s thermal-resilient coherence performance and potential for high-temperature operation.
Fig. 5. : Temperature-dependent coherence of qubit Q1 at charge sweet-spot. a, $T_{1}$ (blue dots, data) versus mixing chamber temperature. The solid curve represents the predicted decay rate $T_{1}(T)=T_{1}(T=0)\cdot\text{tanh}(\hbar\omega_{\rm{q}}/2k_{B}T)$ where we use the measured value at 10 mK for $T_{1}(T=0)$. Extracted electron temperature (gray diamonds, data) versus MXC temperature based on the measured thermal population in the Supplementary. b, Decoherence time, $T_{2}^{*}$ (orange dots) and $T_{2}^{\rm{echo}}$ (green dots) versus the MXC temperature, with power-law fitting. Gray dots and curves show $2T_{1}$ for comparison. c, Extracted pure dephasing time $T_{\phi}$ as a function of the MXC temperature. Red dashed curve: Parameter-free calculation of the resonator-induced dephasing based on Q1 properties.
We then extract the relation between the pure dephasing time ($T_{\phi}$) and MXC temperature, following the coherence decay formula described in the Methods. As shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kc, $T_{\phi}$ begins to decrease significantly when the MXC temperature exceeds 100 mK. In this range, the measured data matches well with a parameter-free model accounting for the effect of resonator thermal photon on the qubit dephasing [24, 50]:
$T_{\phi}^{-1}=\frac{\kappa}{2}\text{Re}\left[\sqrt{\left(1+\frac{2i\chi}{\kappa}\right)^{2}+\frac{8i\chi}{\kappa}n_{\text{th}}}-1\right]$ (3)
where $\kappa$ is the resonator decay rate, and $\chi$ is the resonator’s dispersive shift (see Supplementary IV). $n_{\text{th}}$ represents the resonator thermal photon population, equal to $1/(e^{hf_{\rm{r}}/k_{\text{B}}T}-1)$. At higher temperatures, the degradation of qubit coherence is primarily due to thermal effects. Nevertheless, the deviation of the data from the model below 100 mK suggests the presence of other low-temperature dephasing mechanisms. Separate measurements on Q2 (Supplementary V) give similar results.
Discussion and outlook
In this work, we have quantitatively evaluated, for the first time, the environmental noise isolation provided by the thin neon layer for eNe qubits, demonstrating a number of distinct advantages compared to semiconductor host materials for single electron qubits. Additionally, eNe qubits exhibit thermally resilient coherence, making them promising for high-temperature operation. These systematic coherence and noise characterization highlight the following directions for future research.
The inconsistent spectral properties of eNe qubits indicate that the local neon profile plays a crucial role in the electron-trapping mechanisms. Variations in the neon surface profile, caused by liquid-phase solidification and substrate roughness, introduce randomness in the energy potential landscape that hosts the electron qubits. Additionally, the relative position of trapped electrons to nearby electrodes determines the frequency lever-arm and the qubit’s sensitivity to voltage fluctuations. To enhance consistency and control over individual qubit properties, refined neon growth method and gate-defined electron trapping mechanisms should be developed to mitigate the effects of neon surface randomness.
On the other hand, diverse noise behavior observed in eNe qubits reflects the complexity of the local charge environment. The inconsistent spectral noise distribution suggests locally non-uniform noise sources. Theoretical studies have shown the high sensitivity of qubit’s coherence to fluctuators’ density and distribution [51]. Excess electrons emitted by the tungsten filament could result in artificial charge fluctuators with high mobility on the neon surface. This charge rearrangement can further contribute to the low-frequency qubit drift and occasional electron escape events [52]. Strategies include developing electron loading procedures [29] and protection gates to separate trapped qubits from adjacent charge reservoirs [25] could be implemented to improve the qubit’s stability.
Finally, the limiting factor for the non-radiative energy relaxation rate of eNe qubits needs further investigation. The non-monotonic variations in $T_{1}$ of qubits with different charge sweet-spot frequencies point to localized non-radiative decay channels. The presence of nearby charges may create a sparse bath of two-level fluctuators weakly coupled to the eNe qubit, whose density determines the transverse noise intensity [53, 51]. Electrical manipulations could be applied to bias or repeal those weakly coupled charges to reduce the relaxation rate [54]. With the development of improved electron loading and trapping methods, we foresee the improvement in both coherence and long-term stability of eNe qubits.
Methods
Device and setup
The resonator, electrode, and on-chip filter were patterned with electron beam lithography followed by reactive-ion etching. See details of the TiN film, resonator, and on-chip filter characterization in the Supplementary I. The chip was mounted on a customized printed circuit board within a vacuum-sealed copper cell, which provides direct current (DC) and microwave (MW) interfaces. On the top lid of the cell, a gas filling line was attached to deposit neon at cryogenic temperature, and a tungsten filament was used as the electron source. The cell is mounted on the mixing chamber (MXC) plate of a dilution refrigerator. A total attenuation of 60 dB was applied on the MW input line with infrared filters (QMC-CRYOIRF-004). The MW output line was equipped with cryogenic isolators (LNF-ISISC4_12A) at MXC plate, followed by a high electron mobility transistor (LNF-LNC4_8C) at 4K plate and room temperature amplification. All DC connections were filtered with thermocoaxes cable, LC filters (Mini-Circuits RLP-30+), and homemade low-pass filters with 150 Hz cutoff. Qubit spectroscopic measurements were conducted with a vector network analyzer (Keysight N5222B) and a signal generator (Anritsu MG3692C). Time-domain pulse measurements were conducted with Quantum Machine OPX+ and Octave. The DC gates were applied with QDevil QDAC-II.
Neon growth
Neon is filled with the following procedure. The fridge is warmed up from its base temperature with a heater mounted on the 4 K plate. At this moment, the helium mixture circulation is turned off and all mixture has been collected, while the pulse tube is still on. The heater power is set to a value such that it creates a temperature gradient from 27 K at the 4 K plate to about 25 K at the MXC plate. At such condition, the neon gas is filled and deposit onto the device chip in its liquid phase. After filling, the heater is turned off and whole fridge is cooled down again to let the liquid neon freeze into solid. During the cool down, we further anneal the neon film at 10 K for 1 hour.
Electron deposition
Electrons are ejected from the tungsten filament mounted on the lid of a hermetic copper cell. When the dilution fridge is cooled down to the base temperature, the total resistance on the filament loop is 2 $\Omega$. A current pulse train applied by a pulse generator with $-0.6$ V voltage output, pulse width of 0.1 ms, repetition frequency of 1 kHz, and duration of 0.3 s was used to fire electrons. We noticed that applying higher voltage or longer duration would cause too many electrons to land on the top of the chip, reducing the stability of the trapped electrons.
Noise characterization
We use the dynamical decoupling technique with Carr-Purcell-Meiboom-Gill (CPMG) pulse sequences to study the spectral distribution of high-frequency noise affecting the eNe qubits [37, 38]. In our sequence, $N$ refocusing $Y_{\pi}$ pulses are applied between two $X_{\pi/2}$ pulses with identical separation $\tau/N$ between two pulses. Under a CPMG sequence, qubits’ coherence decaying follows $P_{e}(N,\tau)=P_{0}+a\cdot\text{exp}\left(-\tau/2T_{1}\right)\text{exp}\left(-\chi_{P}\right)\text{exp}\left(-\chi_{N}(\tau)\right)$, accounting for pure dephasing $-\chi_{N}$, energy relaxation $-\tau/2T_{1}$, and decay during driving pulses $-\chi_{P}$ (ref. [37]). We define $T_{2}^{\text{CPMG}}$ as the time when the qubit’s coherence decays by a factor of $1/e$ due to energy relaxation and pure dephasing. Increasing the number of $Y_{\pi}$ pulses reduces the time the qubit is exposed to noise before each refocusing, thereby extending the coherence time. Under Gaussian noise assumption [37, 38], a CPMG sequence with $N$ refocusing pulses and total delay time of $\tau$ imposes a filter function $g_{N}(\omega,\tau)$ to the qubit noise caused by source $\lambda$ with spectral distribution $S_{\lambda}(\omega)$, which determines the qubit’s dephasing $\chi_{N}(\tau)$:
[…]
where $\partial\omega_{\rm{q}}/\partial\lambda$ is the qubit transition frequency’s sensitivity to noise source $\lambda$, the filter function $g_{N}(\omega,\tau)=|y_{N}(\omega,\tau)|^{2}/(\omega\tau)^{2}$, and $y_{N}(\omega,\tau)=1+(-1)^{1+N}\text{exp}(i\omega\tau)$ $+2\sum_{j=1}^{N}(-1)^{j}\text{exp}(i\omega\tau(j-0.5)/N)\text{cos}(\omega\tau_{\pi}/2)$.
With the noise filter imposed by the CPMG sequence, we could approximate the total noise power spectrum density $S$ at frequency $f_{N}$ as (ref. [21, 20]):
$S(2\pi f_{N})=\frac{1}{(T_{\phi})^{2}g_{N}(2\pi f_{N},T_{\phi})\Delta\omega_{N}}$ (5)
where $2\pi f_{N}$ is the peak angular frequency of the first harmonic of $g_{N}(\omega,T_{\phi})$, and $\Delta\omega_{N}$ is the full width at half maximum of the peak.
To investigate the low-frequency dynamics of qubit decoherence near the charge sweet-spot, we repeatedly conducted Ramsey measurements for 128 iterations, with each recording taking 33 seconds [21, 20]. As shown in Fig. Noise-resilient host for electron qubit operation up to 0.4 Kd, we extracted the qubit frequency drift through a fast Fourier transform (FFT) analysis of the measured Ramsey fringes. During the measurement, we also observed more drastic qubit frequency fluctuations on the scale of 10 – 100 MHz, corresponding to 0.14 – 0.44 mV in $\Delta V_{\text{res}}$. Sometimes we even observed the disappear of qubit signatures in the usual gate scanning range. These qubit frequency “jumps” usually occurred less than 0.5 times per hour. We attribute these large fluctuations to charge rearrangements of nearby weakly trapped electrons on neon [52]. For all of presented measurements at various bias points, we ensured that the qubit remained relatively stable and free from such significant fluctuations during data collection.
Acknowledgements.
This work was performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357. D. J., X. H., and X. L. acknowledge support from Argonne National Laboratory Directed Research and Development (LDRD) program. D. J. and Y. W. acknowledge support from the Air Force Office of Scientific Research (AFOSR) under Award No. FA9550-23-1-0636. D. J. and W. G. acknowledge support from the National Science Foundation (NSF) under Award No. OSI-2426768. D. J., X. Zhou, and Y. H. acknowledge support from the Julian Schwinger Foundation for Physics Research. D. J. acknowledges support from the Department of Energy (DOE) under Award No. DE-SC0025542. B. D. acknowledges support from the NSF under Award No. DMR-1906003. C. S. W. and X. L. acknowledge support from Q-NEXT, one of the US DOE Office of Science National Quantum Information Science Research Centers. W. G. acknowledges support from the Gordon and Betty Moore Foundation through Grant DOI 10.37807/gbmf11567 and the National High Magnetic Field Laboratory at Florida State University, which is supported by the NSF Cooperative Agreement No. DMR-2128556 and the state of Florida. X. Zhang and Y. H. acknowledge support from the Office of Naval Research (ONR) Young Investigator Program (YIP) program under Award No. N00014-23-1-2144. X. H. acknowledges support from France and Chicago Collaborating in the Sciences (FACCTS) program. This work was partially supported by the University of Chicago Materials Research Science and Engineering Center, which is funded by the NSF under award number DMR-2011854. This work made use of the Pritzker Nanofabrication Facility of the Institute for Molecular Engineering at the University of Chicago, which receives support from SHyNE, a node of the NSF National Nanotechnology Coordinated Infrastructure (NSF NNCI-1542205). The authors thank David I. Schuster and Amir Yacoby for helpful discussions.
Author contributions X. L., X. Zhou, and D. J. devised the experiment. X. L. and C. S. W. conducted the experiment. X. L. designed and fabricated the device. X. L., C. S. W., and B. D. analyzed the data. Y. H., X. Zhang and X. H. supported the experimental measurement. Y. W. and W. G. supported the theoretical understanding. D. J. conceived the idea and led the project. X. L., C. S. W., X. Zhou, and D. J. wrote the original manuscript. All authors contributed to the work.